1/1

Zhāng Jì,

Maple Bridge Night Mooring

translated by Zhao Yanchun

The moon setting, crows cawing, frost overfilling the skies, The maples riv’rside, lamps aboard, sad the fishermen’s eyes. The toll from the Cold Mountain Temple outside of Suchow Reaches my boat at midnight and floods my sleepless sorrow.

Translation of the ĉina poem “Fēngqiáo Yè Pō” by Zhāng Jì into English by Zhao Yanchun.

Arg-912-1825 (2013-07-27 10:50:16)

http://www.poezio.net/version?poem-id=912&version-id=,1825